# Gauging the Standard Model $1$-form symmetry via gravitational instantons

### v1

@@ 1 Introduction
The Standard Model (SM) of particle physics is believed to be the low-energy effective field theory of some, yet-to-be-constructed, UV-complete theory of gravity. However, in its present form with the known matter content, the SM admits a global electric $\mathbb{Z}_{6}^{(1)}$ 11-form symmetry Tong:2017oea ; Wan:2019gqr ; Anber:2021upc . It is conceivable that this symmetry is only approximate at the energy scales probed so far in terrestrial accelerators, and that as-yet-undiscovered heavy particles could explicitly break it at higher energies Alonso:2024pmq ; Koren:2025utp . Another possibility is that there exists a “desert” between the electroweak and Planck scales, or that any additional heavy degrees of freedom preserve $\mathbb{Z}_{6}^{(1)}$, or at least a nontrivial subgroup thereof.
@@
EH instantons are (anti-)self-dual solutions to the Euclidean Einstein equations with vanishing cosmological constant Eguchi:1978xp ; Eguchi:1978gw . These geometries are said to be asymptotically locally Euclidean (ALE), since instead of approaching the 3-sphere $S^{3}$ at large radial distance $r\rightarrow\infty$ (which is a globally Euclidean space), they asymptotically approach the Lens space $S^{3}/\Gamma$ (only locally Euclidean), where $\Gamma$ is a finite discrete subgroup of $\mathrm{SO}(4)$. In their original formulation, $\Gamma$ was taken $\mathbb{Z}_{2}$, so that the asymptotic boundary becomes $S^{3}/\mathbb{Z}_{2}\equiv\mathbb{RP}^{3}$, the real projective 3-space. The (anti)self-dual ALE metrics are the best understood class of gravitational instantons; see the reviews Eguchi:1980jx ; Dunajski:2024pkf .
@@
Embedding quarks and leptons in the background of EH instantons with $\mathbb{Z}_{6}^{(1)}$ flux supported at the bolt requires additional consistency checks. In particular, one must ensure that the fermions are globally well-defined. A spinor must be continuous along any contractible path connecting two points in the manifold, which imposes nontrivial boundary conditions at the bolt in the presence of gauge fields. These boundary conditions encode topological data that is transmitted to the asymptotic boundary $\mathbb{RP}^{3}$. For a Dirac spinor $\Psi$ in the presence of a $U(1)$ gauge field, it is sufficient to impose the boundary condition at the bolt:
$\Psi(r=a,\theta,\varphi,\psi+2\pi)=(-1)^{\cal C}\,\Psi(r=a,\theta,\varphi,\psi)\,.$
@@
The identification of the asymptotic region under the discrete $\mathbb{Z}_{2}$ symmetry, which gives rise to the boundary $\mathbb{RP}^{3}$, implies that the space surrounding the EH instanton is effectively a half-space. In this framework, the EH instanton can be interpreted as mediating the transition amplitude between a specially prepared initial state entangled between the two identical halves of the space (left $L$ and right $R$; see Footnote 17 for a discretized version of such state):
@@ 2.1 Turning on a $U(1)$ gauge field
$F_{(1)}=dA_{U(1)}=\frac{{\cal C}a^{2}}{r^{3}}dr\wedge(d\psi+\cos\theta d\varphi)+\frac{{\cal C}a^{2}}{2r^{2}}\sin\theta d\varphi\wedge d\theta$ (9)
@@
It follows that the topological charge of the $U(1)$ field is [^6] It also helps to compare this result with the topological charge of a $U(1)$ gauge field in the background of the complex projective space $\mathbb{CP}^{2}$, a nonspin manifold. Here, the space admits a normalizable harmonic $2$-form $\cal K$} such that the integral of $\cal K$} over the $2$-cycle, in this case it is $\mathbb{CP}^{1}\subset\mathbb{CP}^{2}$, is $\int_{\mathbb{CP}^{1}}{\cal K}=1$. The self-intersection number is given by the wedge product $\int_{\mathbb{CP}^{2}}{\cal K}\wedge{\cal K}=1$. A self-dual $U(1)$ gauge field is given by $F_{(1)}=-2\pi\,{\cal C}\,{\cal K}$, and thus, its associated topological charge is $Q_{(1)}^{\mathbb{CP}^{2}}=\frac{C^{2}}{2}\,.$ (17) See, e.g., the appendices of Anber:2020gig for the details of the calculations. We see that the different self-intersection numbers of EH and $\mathbb{CP}^{2}$ spaces explain the extra factor of $1/2$ that multiplies the topological charge in EH space.
@@ 2.3 Spinors in EH space in the presence of abelian fields
Consider the parallel transport of a Dirac fermion $\Psi$, carrying a positive unit electric charge under $U(1)$, along the contractible path $\ell_{r=a}$ connecting $(r=a,\theta=0,\varphi=0,\psi)$ and $(r=a,\theta=0,\varphi=0,\psi+2\pi)$ (or any path at $r=a$ with constant $\theta,\varphi$) on the bolt $S^{2}$. Along this path, the fermion acquires both $U(1)$ and gravitational holonomies:
$\Psi(r,\theta,\varphi,\psi+2\pi)$ $=\exp\left[i\int_{\ell_{r=a}}A\right]\,\exp\left[-\tfrac{1}{2}\int_{\ell_{r=a}}\omega_{ab}\Sigma^{ab}\right]\Psi(r,\theta,\varphi,\psi)\,$, (24)
@@
$\exp\left[-\tfrac{1}{2}\int_{\ell_{r=a}}\omega_{ab}\Sigma^{ab}\right]=I_{4}\,,\quad\exp\left[i\int_{\ell_{r=a}}A\right]=e^{i\pi\mathcal{C}}.$ (25)
Thus, the gravitational holonomy is trivial along $\ell_{r=a}$, reflecting the fact that EH spaces are spin. Only the gauge holonomy is generally nontrivial, so that
$\text{at }r=a:\qquad\Psi(r,\theta,\varphi,\psi+2\pi)=(-1)^{\cal C}\,\Psi(r,\theta,\varphi,\psi)\,.$ (26)
@@
The computation of the gauge holonomy $e^{i\int_{\ell_{\infty}}A}$ requires more care. Naively, because $A$ is rapidly decaying, one would conclude that this holonomy is trivial, $(+1)$. This, however, is not correct. The idea is that such a quick conclusion overlooks the possibility that topological data may be transmitted from the interior to the boundary, which is indeed our case. To correctly compute the gauge holonomy, we use Stokes’ theorem $\int_{\ell_{\infty}}A=\int_{\Sigma}F$, where $\Sigma$ is a $2$-surface that interpolates between the loop $\ell_{\infty}$ on the boundary $\mathbb{RP}^{3}$ and the bolt in the interior. Topologically, it is a cigar shape stretching from the boundary, capping off at the bolt (the $\psi$-circle shrinks smoothly at the bolt [^8] See the EH metric (132) in the limit $r\rightarrow a$, which helps with the visualization.). Using the expression of $F_{(1)}$ in (9), fixing $\theta,\varphi$, and integrating over $0\leq\psi\leq 2\pi$, $a\leq r\leq\infty$, we readily find
@@
Before concluding this section, let us also note that a scalar $h$, which carries a unit electric charge under $U(1)$, acquires a gauge holonomy upon parallel transport along a contractible path at the bolt. Similarly, the holonomy of the flat line bundle at infinity picks a (1)(-1) sign when $\cal C$} is odd. Thus, the boundary conditions for the scalar are:
$\text{at }r=a:\qquad h(r,\theta,\varphi,\psi+2\pi)=(-1)^{\cal C}\,h(r,\theta,\varphi,\psi)\,$,
@@ 2.4 Dirac fermion zero modes in EH space
The ket $|j,m^{\prime},m\rangle$ is the Wigner-D matrix Sakurai:2011zz , which is the generalization of spherical harmonics on $\mathbb{S}^{3}$. Here, $j$ is the total angular momentum, which can be a positive half-integer, positive integer, or 0: $j=0,\frac{1}{2},1,$.... The values of $m^{\prime},m$ are in the range $|m|,|m^{\prime}|\leq j$. The functions $g_{1},g_{2}$ atisfy the first-order differential equations
@@ 3 Background of $U(N)$ gauge bundle and fermion zero modes
This obstruction can be understood in several ways. For instance, consider a fundamental fermion subject to the holonomy of the gauge field given in (19). Repeating the analysis starting from (24), we find that upon traversing a contractible path at the bolt, the fermion acquires the phase
$\Psi(r=a,\theta,\varphi,\psi+2\pi)$ $=$ $\exp\!\left[i\pi\,m_{(N)}\,\mathrm{diag}_{N\times N}\!\left(1-\tfrac{1}{N},\,-\tfrac{1}{N},\,\dots,\,-\tfrac{1}{N}\right)\right]$ (43)
$\times\Psi(r=a,\theta,\varphi,\psi)\,.$
@@
In order for the $SU(N)$ spinor to be globally well-defined under parallel transport along a contractible loop $\psi\to\psi+2\pi$ on $S^{2}$, one needs to restrict the total angular momentum $j$ of each component of the $N\times 1$ coloumn vector $\lambda_{2}$ to either an integer or half-integer according to whether the enteries in $\mbox{diag}_{N\times N}\!\left(m_{(N)}+{\cal C},{\cal C},\dots,{\cal C}\right)$ are even or odd. Thus, we have $4$ distinct cases
@@
$=$ $\mbox{diag}_{N\times N}\!\left((-1)^{m_{(N)}+{\cal C}},(-1)^{{\cal C}},\dots,(-1)^{{\cal C}}\right)$
@@
Now, we have all the machinery needed to apply our construction to the Standard Model (SM) when placed in the background of EH sapce.
@@ 4.1 Background of $\mathbb{Z}_{3}^{(1)}$ 11-form symmetry
Inspecting (4), we remark that transporting any $SU(2)$-charged field—namely $q_{L}$, $l_{L}$, and hh—around a closed loop at the bolt results in a trivial or phase of $\pi$ (holonomy of (1)(-1)), depending on whether $\cal C$} is even or odd. In contrast, all fields uncharged under $SU(2)$ invariably acquire a trivial phase. This distinction originates from the fact that all $SU(2)$-charged fields carry odd charges under the $U(1)$ hypercharge gauge symmetry. Therefore, in applying the APS index theorem, and whenever $\cal C$} is even, one must always select the boundary flat line bundle with a trivial holonomy that yields $\eta=-1/8$ in (C). However, when $\cal C$} is odd, the sign of the $\eta$-invariant will depend on the fermion species: we must take $\eta=+1/8$ in (C) for fields charged under $SU(2)$ (all such fields carry odd charges under $U(1)$ hypercharge gauge field) and $\eta=-1/8$ for the rest of fields. We remind the reader that the two possible twists, $\{+1,-1\}$, on $\mathbb{RP}^{3}$ arise from the bulk background gauge field. When restricted to the boundary, this bulk field defines a flat line bundle whose holonomy determines the choice of twist, and hence fixes one of the two possible values of the $\eta$-invariant.
@@
The resulting vertex mediates baryon- and lepton-number violating processes, with the net violation (per family) given by
@@ 5 EH spaces as saddles in the path integral: gauging the SM $1$-form symmetry
$A_{\mu}$ are the fluctuations around the EH background $A_{\mu}^{(B)}$, and both respect the antipodal constraint [^14] In the EH geometry, the anti-podal identification is $(x,y,z,t)\;\sim\;(-x,-y,-z,-t)$. In our path integral treatment, however, it is sufficient to impose a constraint on the initial data at some Euclidean time slice $t=t_{0}$. Specifically, for the gauge potential, we require $A_{\mu}(x,y,z,t_{0})\;=\;-A_{\mu}(-x,-y,-z,t_{0})$, which enforces the correct parity under spatial inversion. The Euclidean time evolution then propagates this condition consistently into the bulk.. We shall assume $t_{0}>a$, since the region $x^{2}+y^{2}+z^{2}+t^{2}<a^{2}$ is excluded [^15] This, however, does not imply that the spacetime ends abruptly at the bolt; EH space is smooth and geodesically complete, and elliptic self-adjoint operators in this space, e.g., the Laplacian, the Dirac operator, etc., furnish a complete set of eigenstates without imposing further boundary conditions at “the would be end of space”.. Condition (95) reflects the fact that the gauge bundle on EH has a nontrivial $\mathbb{Z}_{2}$ transition function (of geometric origin) across the antipodal map, which can be made explicit in the chosen gauge. This constraint is not arbitrary; it is a physical requirement, up to a gauge transformation, if we want a path integral over $A_{\mu}$ to describe the transition amplitude in the background of the EH instanton [^16] To add to Footnote 14, one could in principle choose the initial slice at $t_{0}=0$, in which case the antipodal identification would imply $A_{\mu}(x,y,z,0)\;=\;-\,A_{\mu}(-x,-y,-z,0)$. However, the state defined at $t_{0}=0$, namely $|\mathrm{EH}\rangle_{0}$, is not well defined because the geometry requires $x^{2}+y^{2}+z^{2}+t^{2}\geq a^{2}$ (see tHooft:1988wxy for further discussion). The resolution is to evolve this formal state forward in Euclidean time using the Hamiltonian $H$, leading to the well-defined state $|\mathrm{EH}\rangle_{t_{0}}\;=\;e^{-t_{0}H}\,|\mathrm{EH}\rangle_{0}$ ..
@@
To ensure that the path integral respects the unitarity and locality (cluster decomposition), one must also sum over all integer topological sectors [^18] This assertion stems from our experience with Yang-Mills theory, see, e.g.,Weinberg:1996kr ; Donnelly:2013tia ; Tanizaki:2019rbk ; Anber:2025vjo .. For the $U(1)$ hypercharge gauge field, this entails a sum over $\cal C$$\in\mathbb{Z}$ (as well as sums over the integer topological charges of the $SU(3)\times SU(2)$ sectors, which we do not discuss here). Thus, in a fixed $m_{(3)},m_{(2)}$ background, summing over $\cal C$}, and gauge-fixing via the Fadeev–Popov procedure, the physical amplitude takes the form
@@
Let us emphasize an important point in the above setup. In writing (101, 102), we assumed a fixed EH background and included a sum over all $\mathbb{Z}_{6}^{(1)}$ flux sectors supported by the bolt. It is useful to contrast this with the case of $U(1)$ gauge theory: there, summing over all integer flux sectors $\cal C$$\in\mathbb{Z}$ is mandatory, since otherwise the path integral would violate unitarity and cluster decomposition. By comparison, summing over distinct $(m_{(3)},m_{(2)})$ is a matter of physical choice: performing the sum corresponds to gauging the discrete $\mathbb{Z}_{6}^{(1)}$ one-form symmetry. If one does not perform this sum, the theory is defined in a fixed background of the $\mathbb{Z}_{6}^{(1)}$ symmetry; if one does perform the sum, the symmetry is dynamically gauged. Thus, the correct procedure for computing physical observables (e.g., correlation functions or condensates) is as follows: first, compute the path integral in a given topological sector, then
  • always sum over $\cal C$$\in\mathbb{Z}$ to enforce unitarity and cluster decomposition in the semi-classical sense; the latter ensures that widely separated operators approximately factorize at large distances. We emphasize that in full quantum gravity, cluster decomposition may not be strictly valid due to effects such as topology change or wormholes, but it provides a guiding principle in the semi-classical regime;

### v2

@@ 1 Introduction
The Standard Model (SM) of particle physics is believed to be the low-energy effective field theory of some, yet-to-be-constructed, UV-complete theory of gravity. However, in its present form with the known matter content, the SM admits a global electric $\mathbb{Z}_{6}^{(1)}$ 11-form symmetry Tong:2017oea ; Wan:2019gqr ; Anber:2021upc . It is conceivable that this symmetry is only approximate at the energy scales probed so far in terrestrial accelerators, and that as-yet-undiscovered heavy particles could explicitly break it at higher energies Alonso:2024pmq ; Li:2024nuo ; Koren:2025utp . Another possibility is that there exists a “desert” between the electroweak and Planck scales, or that any additional heavy degrees of freedom preserve $\mathbb{Z}_{6}^{(1)}$, or at least a nontrivial subgroup thereof.
@@
EH instantons are (anti-)self-dual solutions to the Euclidean Einstein equations with vanishing cosmological constant Eguchi:1978xp ; Eguchi:1978gw . These geometries are said to be asymptotically locally Euclidean (ALE), since instead of approaching the 3-sphere $S^{3}$ at large radial distance $r\rightarrow\infty$ (which is a globally Euclidean space), they asymptotically approach the Lens space $S^{3}/\Gamma$ (only locally Euclidean), where $\Gamma$ is a finite discrete subgroup of $SU(2)$. In their original formulation, $\Gamma$ was taken $\mathbb{Z}_{2}$, so that the asymptotic boundary becomes $S^{3}/\mathbb{Z}_{2}\equiv\mathbb{RP}^{3}$, the real projective 3-space. The (anti)self-dual ALE metrics are the best understood class of gravitational instantons; see the reviews Eguchi:1980jx ; Dunajski:2024pkf .
@@
Embedding quarks and leptons in the background of EH instantons with $\mathbb{Z}_{6}^{(1)}$ flux supported at the bolt requires additional consistency checks. In particular, one must ensure that the fermions are globally well-defined. A spinor must be continuous along any contractible path connecting two points in the manifold, which imposes nontrivial boundary conditions near the bolt in the presence of gauge fields. These boundary conditions encode topological data that is transmitted to the asymptotic boundary $\mathbb{RP}^{3}$. For a Dirac spinor $\Psi$ in the presence of a $U(1)$ gauge field, it is sufficient to impose the boundary condition near the bolt:
$\Psi(r\approx a,\theta,\varphi,\psi+2\pi)=(-1)^{\cal C}\,\Psi(r\approx a,\theta,\varphi,\psi)\,.$
@@
The identification of the asymptotic region under the discrete $\mathbb{Z}_{2}$ symmetry, which gives rise to the boundary $\mathbb{RP}^{3}$, implies that the space surrounding the EH instanton is effectively a half-space. In this framework, the EH instanton can be interpreted as mediating the transition amplitude between a specially prepared initial state entangled between the two identical halves of the space (left $L$ and right $R$; see Footnote 19 for a discretized version of such state):
@@
Although the EH space is asymptotic to $\mathbb{R}^{4}/\mathbb{Z}_{2}$ rather than strictly $\mathbb{R}^{4}$, its asymptotic region is locally indistinguishable from flat space: on scales small compared to the curvature radius, and within a simply connected neighbourhood that does not wind the nontrivial $1$-cycle, the induced metric and SM background fields coincide with those of $\mathbb{R}^{4}$. Thus no local observer confined to such a region can distinguish $S^{3}$ from $S^{3}/\mathbb{Z}_{2}$, and ordinary correlators of point operators with all insertions in that region agree with those in flat space. The difference lies only in the global topology, which, once the SM $1$-form symmetry is gauged, modifies the allowed topological sectors of the gauge bundle and thereby the structure of instanton-induced multi-fermion operators. In this sense EH space is a perfectly admissible Euclidean saddle contributing to the path integral with asymptotically locally flat (ALE) boundary conditions. In fact, EH space is the minimal and most analytically tractable ALE space, and thus can be viewed as a representative of a broader class of configurations that, outside a compact region, locally coincide with flat $\mathbb{R}^{4}$ but have nontrivial topology in the interior.
@@ 2.1 Turning on a $U(1)$ gauge field
$F_{(1)}=dA_{U(1)}=-\frac{{\cal C}a^{2}}{r^{3}}dr\wedge(d\psi+\cos\theta d\varphi)+\frac{{\cal C}a^{2}}{2r^{2}}\sin\theta d\varphi\wedge d\theta$ (9)
@@
It follows that the topological charge of the $U(1)$ field is [^6] It also helps to compare this result with the topological charge of a $U(1)$ gauge field in the background of the complex projective space $\mathbb{CP}^{2}$, a nonspin manifold. Here, the space admits a normalizable harmonic $2$-form $\cal K$} such that the integral of $\cal K$} over the $2$-cycle, in this case it is $\mathbb{CP}^{1}\subset\mathbb{CP}^{2}$, is $\int_{\mathbb{CP}^{1}}{\cal K}=1$. The self-intersection number is given by the wedge product $\int_{\mathbb{CP}^{2}}{\cal K}\wedge{\cal K}=1$. A self-dual $U(1)$ gauge field is given by $F_{(1)}=-2\pi\,{\cal C}\,{\cal K}$, and thus, its associated topological charge is $Q_{(1)}^{\mathbb{CP}^{2}}=\frac{{\cal C}^{2}}{2}\,.$ (17) See, e.g., the appendices of Anber:2020gig for the details of the calculations. We see that the different self-intersection numbers of EH and $\mathbb{CP}^{2}$ spaces explain the extra factor of $1/2$ that multiplies the topological charge in EH space.
@@ 2.3 Spinors in EH space in the presence of abelian fields
Consider the parallel transport of a Dirac fermion $\Psi$, carrying a positive unit electric charge under $U(1)$, along the contractible path $\ell_{r\approx a}$ connecting $(r\approx a,\theta=0,\varphi=0,\psi)$ and $(r\approx a,\theta=0,\varphi=0,\psi+2\pi)$ (or any path with constant $\theta,\varphi$) near the bolt [^8] We evaluate the holonomy along a small contractible loop $\ell_{r=a+\epsilon}$ and then take the limit $\varepsilon\to 0$. In our coordinates, the frame becomes singular exactly at $r=a$ (i.e. the metric degenerates to $S^{2}$), so it is technically more convenient to work ”near the bolt” and understand all conditions as limits $r\to a$ from above. $S^{2}$ (the metric near $r\approx a$ is given by (132)). Along this path, the fermion acquires both $U(1)$ and gravitational holonomies:
$\Psi(r,\theta,\varphi,\psi+2\pi)$ $=\exp\left[i\int_{\ell_{r\approx a}}A\right]\,\exp\left[-\tfrac{1}{2}\int_{\ell_{r\approx a}}\omega_{ab}\Sigma^{ab}\right]\Psi(r,\theta,\varphi,\psi)\,$, (24)
@@
$\exp\left[-\tfrac{1}{2}\int_{\ell_{r\approx a}}\omega_{ab}\Sigma^{ab}\right]=I_{4}\,,\quad\exp\left[i\int_{\ell_{r\approx a}}A\right]=e^{i\pi\mathcal{C}}.$ (25)
Thus, the gravitational holonomy is trivial along $\ell_{r\approx a}$, reflecting the fact that EH spaces are spin. Only the gauge holonomy is generally nontrivial, so that
$\text{at }r\approx a:\qquad\Psi(r,\theta,\varphi,\psi+2\pi)=(-1)^{\cal C}\,\Psi(r,\theta,\varphi,\psi)\,.$ (26)
@@
Notice here that, unlike the case of parallel transporting near the bolt, the gravitational holonomy is non-trivial. The computation of the gauge holonomy $e^{i\int_{\ell_{\infty}}A}$ requires more care. Naively, because $A$ is rapidly decaying, one would conclude that this holonomy is trivial, $(+1)$. This, however, is not correct. The idea is that such a quick conclusion overlooks the possibility that topological data may be transmitted from the interior to the boundary, which is indeed our case. To correctly compute the gauge holonomy, we use Stokes’ theorem $\int_{\ell_{\infty}}A=\int_{\Sigma}F$, where $\Sigma$ is a $2$-surface that interpolates between the loop $\ell_{\infty}$ on the boundary $\mathbb{RP}^{3}$ and the bolt in the interior. Topologically, it is a cigar shape stretching from the boundary, capping off at the bolt (the $\psi$-circle shrinks smoothly at the bolt [^9] See the EH metric (132) in the limit $r\rightarrow a$, which helps with the visualization.). Using the expression of $F_{(1)}$ in (9), fixing $\theta,\varphi$, and integrating over $0\leq\psi\leq 2\pi$, $a\leq r\leq\infty$, we readily find
@@
Before concluding this section, let us also note that a scalar $h$, which carries a unit electric charge under $U(1)$, acquires a gauge holonomy upon parallel transport along a contractible path near the bolt. Similarly, the holonomy of the flat line bundle at infinity picks a (1)(-1) sign when $\cal C$} is odd. Thus, the boundary conditions for the scalar are:
$\text{at }r\approx a:\qquad h(r,\theta,\varphi,\psi+2\pi)=(-1)^{\cal C}\,h(r,\theta,\varphi,\psi)\,$,
@@ 2.4 Dirac fermion zero modes in EH space
The ket $|j,m^{\prime},m\rangle$ is the Wigner-D matrix Sakurai:2011zz , which is the generalization of spherical harmonics on $\mathbb{S}^{3}$. Here, $j$ is the total angular momentum, which can be a positive half-integer, positive integer, or 0: $j=0,\frac{1}{2},1,$.... The values of $m^{\prime},m$ are in the range $|m|,|m^{\prime}|\leq j$. The functions $g_{1},g_{2}$ satisfy the first-order differential equations
@@ 3 Background of $U(N)$ gauge bundle and fermion zero modes
This obstruction can be understood in several ways. For instance, consider a fundamental fermion subject to the holonomy of the gauge field given in (19). Repeating the analysis starting from (24), we find that upon traversing a contractible path near the bolt, the fermion acquires the phase
$\Psi(r\approx a,\theta,\varphi,\psi+2\pi)$ $=$ $\exp\!\left[i\pi\,m_{(N)}\,\mathrm{diag}_{N\times N}\!\left(1-\tfrac{1}{N},\,-\tfrac{1}{N},\,\dots,\,-\tfrac{1}{N}\right)\right]$ (43)
$\times\Psi(r\approx a,\theta,\varphi,\psi)\,.$
@@
In order for the $SU(N)$ spinor to be globally well-defined under parallel transport along a contractible loop $\psi\to\psi+2\pi$ on $S^{2}$, one needs to restrict the total angular momentum $j$ of each component of the $N\times 1$ column vector $\lambda_{2}$ to either an integer or half-integer according to whether the entries in $\mbox{diag}_{N\times N}\!\left(m_{(N)}+{\cal C},{\cal C},\dots,{\cal C}\right)$ are even or odd. Thus, we have $4$ distinct cases
@@
$=$ $-\mbox{diag}_{N\times N}\!\left((-1)^{m_{(N)}+{\cal C}},(-1)^{{\cal C}},\dots,(-1)^{{\cal C}}\right)$
@@
Now, we have all the machinery needed to apply our construction to the Standard Model (SM) when placed in the background of EH space.
@@ 4.1 Background of $\mathbb{Z}_{3}^{(1)}$ 11-form symmetry
Inspecting (4), we remark that transporting any $SU(2)$-charged field—namely $q_{L}$, $l_{L}$, and hh—around a closed loop near the bolt results in a trivial or phase of $\pi$ (holonomy of (1)(-1)), depending on whether $\cal C$} is even or odd. In contrast, all fields uncharged under $SU(2)$ invariably acquire a trivial phase. This distinction originates from the fact that all $SU(2)$-charged fields carry odd charges under the $U(1)$ hypercharge gauge symmetry. Therefore, in applying the APS index theorem, and whenever $\cal C$} is even, one must always select the boundary flat line bundle with a trivial holonomy that yields $\eta=-1/8$ in (C). However, when $\cal C$} is odd, the sign of the $\eta$-invariant will depend on the fermion species: we must take $\eta=+1/8$ in (C) for fields charged under $SU(2)$ (all such fields carry odd charges under $U(1)$ hypercharge gauge field) and $\eta=-1/8$ for the rest of fields. We remind the reader that the two possible twists, $\{+1,-1\}$, on $\mathbb{RP}^{3}$ arise from the bulk background gauge field. When restricted to the boundary, this bulk field defines a flat line bundle whose holonomy determines the choice of twist, and hence fixes one of the two possible values of the $\eta$-invariant.
@@
The resulting vertex mediates baryon- and lepton-number violating processes, with the net violation (per family) given by [^14] Discussions of CP violation in QED in the background of EH instantons appeared before in Arunasalam:2018eaz ; see also Deser:1980kc .
@@ 5 EH spaces as saddles in the path integral: gauging the SM $1$-form symmetry
$A_{\mu}$ are the fluctuations around the EH background $A_{\mu}^{(B)}$, and both respect the antipodal constraint [^16] In the EH geometry, the anti-podal identification is $(x,y,z,t)\;\sim\;(-x,-y,-z,-t)$. In our path integral treatment, however, it is sufficient to impose a constraint on the initial data at some Euclidean time slice $t=t_{0}$. Specifically, for the gauge potential, we require $A_{\mu}(x,y,z,t_{0})\;=\;-A_{\mu}(-x,-y,-z,t_{0})$, which enforces the correct parity under spatial inversion. The Euclidean time evolution then propagates this condition consistently into the bulk.. We shall assume $t_{0}>a$, since the region $x^{2}+y^{2}+z^{2}+t^{2}<a^{2}$ is excluded [^17] This, however, does not imply that the spacetime ends abruptly at the bolt; EH space is smooth and geodesically complete, and elliptic self-adjoint operators in this space, e.g., the Laplacian, the Dirac operator, etc., furnish a complete set of eigenstates without imposing further boundary conditions at “the would be end of space”.. Condition (95) reflects the fact that the gauge bundle on EH has a nontrivial $\mathbb{Z}_{2}$ transition function (of geometric origin) across the antipodal map, which can be made explicit in the chosen gauge. This constraint is not arbitrary; it is a physical requirement, up to a gauge transformation, if we want a path integral over $A_{\mu}$ to describe the transition amplitude in the background of the EH instanton [^18] To add to Footnote 16, one could in principle choose the initial slice at $t_{0}=0$, in which case the antipodal identification would imply $A_{\mu}(x,y,z,0)\;=\;-\,A_{\mu}(-x,-y,-z,0)$. However, the state defined at $t_{0}=0$, namely $|\mathrm{EH}\rangle_{0}$, is not well defined because the geometry requires $x^{2}+y^{2}+z^{2}+t^{2}\geq a^{2}$ (see tHooft:1988wxy for further discussion). The resolution is to evolve this formal state forward in Euclidean time using the Hamiltonian $H$, leading to the well-defined state $|\mathrm{EH}\rangle_{t_{0}}\;=\;e^{-t_{0}H}\,|\mathrm{EH}\rangle_{0}$ ..
@@
To ensure that the path integral respects locality (cluster decomposition), one must also sum over all integer topological sectors [^20] This assertion stems from our experience with Yang-Mills theory, see, e.g.,Weinberg:1996kr ; Donnelly:2013tia ; Tanizaki:2019rbk ; Anber:2025vjo .. For the $U(1)$ hypercharge gauge field, this entails a sum over $\cal C$$\in\mathbb{Z}$ (as well as sums over the integer topological charges of the $SU(3)\times SU(2)$ sectors, which we do not discuss here). Thus, in a fixed $m_{(3)},m_{(2)}$ background, summing over $\cal C$}, and gauge-fixing via the Fadeev–Popov procedure, the physical amplitude takes the form
@@
Let us emphasize an important point in the above setup. In writing (101, 102), we assumed a fixed EH background and included a sum over all $\mathbb{Z}_{6}^{(1)}$ flux sectors supported by the bolt. It is useful to contrast this with the case of $U(1)$ gauge theory: there, summing over all integer flux sectors $\cal C$$\in\mathbb{Z}$ is mandatory, since otherwise the path integral would violate cluster decomposition. By comparison, summing over distinct $(m_{(3)},m_{(2)})$ is a matter of physical choice: performing the sum corresponds to gauging the discrete $\mathbb{Z}_{6}^{(1)}$ one-form symmetry. If one does not perform this sum, the theory is defined in a fixed background of the $\mathbb{Z}_{6}^{(1)}$ symmetry; if one does perform the sum, the symmetry is dynamically gauged. Thus, the correct procedure for computing physical observables (e.g., correlation functions or condensates) is as follows: first, compute the path integral in a given topological sector, then
  • always sum over $\cal C$$\in\mathbb{Z}$ to enforce cluster decomposition in the semi-classical sense; the latter ensures that widely separated operators approximately factorize at large distances. We emphasize that in full quantum gravity, cluster decomposition may not be strictly valid due to effects such as topology change or wormholes, but it provides a guiding principle in the semi-classical regime;